# Frostman random variables, entropy inequalities, and applications

### v3

@@ 2.4 Frostman measures
(iii) We say a random variable $X$ is $(s,C)$-Frostman given $Y$, if
@@
for any ball $B(x,r)$ and Borel set $E$. Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
The following properties follows directly from the definitions, whose proofs are left as an exercise.
@@
(iv) Suppose that $(X,Y)$ is a jointly $(s_{1},s_{2},C)$-Frostman random variable whose range is contained in $B(0,R)$, where $R>0$. Then $(X,Y)$ is marginally $(s_{1},CR^{s_{2}};s_{2},CR^{s_{1}})$-Frostman.
@@
Assume that $b\neq 0$. If $q_{i}>0$, then $Y$ is located in an interval $J_{i}$ of length at most $\frac{2(|a|+1)r}{|b|}$. So
@@
First, let us consider the case $ad\neq 0$. When $r_{1}\leq r_{2}$, let $\{J_{i}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r_{1}$. We estimate $p$ by
@@
One can check that, if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{2(|b|+1)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
@@
We apply similar arguments as above. For non-zero terms, the value of $Y$ locates in an interval $J_{i}^{\prime}$ of length at most $\frac{2(|c|+1)r_{2}}{|d|}$. So
@@
(i) is taken from [12, Lemma 4.8].
@@
Let $m\geq 1$. Suppose that $Z$ is a random variable with bounded range in ${\mathbb{R}}^{m}$, and $f$ is a function such that $domain(f)\supseteq range_{+}(Z)$. Assume that there are numbers $s,C>0$ such that $P(|f(Z)|<r)\leq Cr^{s}$ for any $r>0$. Let $0<\eta\leq\min\{2^{-1}(2C)^{-1/s},\,1/2\}$, and denote by $E$ the event that $|Z|\geq\eta$. Assume that $P(E)\gg 1$. Then the random variable $f(Z)_{E}$ satisfies
$H_{n}(f(Z)_{E})=\frac{H_{n}(f(Z))}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
@@
Recall that $f(Z)_{E}$ is the random variable such that $f(Z)_{E}\equiv(f(Z)|E)$, and $P_{E}$ is the conditional probability measure given $E$ (see Lemma A.2).
Assume that $k_{0}$ is a natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$. Combining the conditions in this lemma, one deduces that
$H_{n}(f(Z)_{E})$ $=-\sum\limits_{k\in\mathbb{Z}^{m}\atop|k|\geq k_{0}+m}P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\log P_{E}(D_{n}(f(Z)_{E})=k2^{-n})$
$\qquad+O\left(-\sum\limits_{k_{0}-m\leq|k|<k_{0}+m}P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\log P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\right)$
$=-\sum\limits_{|k|\geq k_{0}+m}\frac{P(D_{n}(f(Z))=k2^{-n})}{P(E)}\big{(}\log P(D_{n}(f(Z))=k2^{-n})-\log P(E)\big{)}+O_{m}(1)$
$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})+O_{m}(1).$
First, let us consider the case that $\eta>m2^{-n+2}$, which gives $(k_{0}+2m)2^{-n}\leq\eta+(2m+1)2^{-n}<2\eta$. Applying Lemma 2.9 with the fact
$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\leq P(|f(Z)|<(k_{0}+2m)2^{-n})\leq P(|f(Z)|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
@@
$H(f(Z)_{E})$ $=\frac{1}{P(E)}\left(H_{n}(f(Z))+\sum\limits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})\right)+O_{m}(1)$
$=\frac{H_{n}(f(Z))}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
@@ 2.5 Entropy of the distance function
Let $\mu$ be a probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, for some $R>0$. Then, for any $0<t\leq 1$ and for real numbers $0<\eta<r<2R$, we have $\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)$.
@@
Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$ and $C,R>0$. Then for any $r,\eta,t$ with $0<\eta<r<2R$, $0<t\leq 1$ and $t<2s-\frac{1}{2}$, it satisfies that
@@
$\qquad=O_{t}\left(\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}\right)+O_{t}(1):=M.$
@@
$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(\eta^{t}).$
@@
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and s<t<2s1/2s<t<2s-1/2, we have
@@
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
@@
By Lemma 2.34, one has $P(|Z|<r)\leq Cr^{2s}$ for any $r>0$. Let $\eta$ be a parameter with $0<\eta<\min\{2^{-1}(2C)^{-1/2s},\,1/2\}$ to be determined later. Denote by $E$ the event that $|Z|\geq\eta$. By Lemma 2.33, we deduce that
@@
$H_{n}(Z^{2})$ $\geq P(E)\,H_{n}(Z_{E}^{2})\geq P(E)\left(H_{n}(Z_{E})+O\left(\log\frac{1}{\eta}\right)+O(1)\right)$
@@
One can verify that the requirements involving $t$ and $\eta$ are satisfied, and $2s-t-1/2\gg_{s}1$. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
@@
Let $X$, $X^{\prime}$, $Y$, $Y^{\prime}$ be independent real random variables. Suppose that $X$ and $X^{\prime}$ are both $(s,C)$-Frostman with ranges contained in [c,c][-c,c], where $0<s\leq 1$ and $C,c>0$. Then, for any $n\in\mathbb{N}$,
@@
By Lemma 2.31, the random variable $Z:=X-X^{\prime}$ is $(s,O(C))$-Frostman. By Lemma 2.32, one has $H_{n}(Z)\geq ns-\log C+O(1)$. Also, $range(Z)\subseteq[-2c,2c]$. Let $E$ be the event that $|Z|\geq\eta$ as in Lemma 2.33. Follow the same arguments as in the proof of Theorem 2.41, we obtain that
@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
In this section, we follow the method introduced by Tao in [26, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem. We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section. The Balog-Szemerédi-Gowers theorem (or one of its many variants) says that if one has a finite subset of the integers whose sums line up a lot, then it must not have many sums. More precisely, given $A\subseteq\mathbb{Z},$ whose elements $a_{j}\in A$ have $\gamma|A|^{3}$ relations of the form $a_{1}+a_{2}=a_{3}+a_{4},$ there must be a large subset $A^{\prime}\subseteq A,$ where $|A^{\prime}|\geq C_{1}(\gamma)|A|$ that satisfies $|A^{\prime}+A^{\prime}|\leq C_{2}(\gamma)|A^{\prime}|,$ where the constants $C_{j}$ above are positive and only depend on $\gamma.$
@@
Let $n$ be a natural number, and $(X,Y)$ be a given random variable. Suppose that
@@
In addition, for any weakly $\kappa$-bi-Lipschitz functions $f$ and $g$ with
@@
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa).$ (28)
@@
Let $n$ be a natural number, and $(X,Y)$ be a given random variable. Then there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that (i)–(iv) in Theorem 3.1 hold, with $X^{\prime\prime}$ and $Y^{\prime\prime}$ being conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$.
@@
Let $n$ be a natural number and $(X,Y)$ be a given random variable. Let $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ be the random variable shown in Lemma 3.3. For any $\kappa$-Lipschitz function $h$, the following properties hold.
@@
Noting that $(Y^{\prime},D_{n}(X^{\prime}))\equiv(Y^{\prime\prime},D_{n}(X^{\prime}))$, one asserts that
@@
Combining this with (i), one gets (ii). Claim (iii) follows by a similar argument. ∎
@@
With $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ and $f,g$ as stated in Theorem 3.1, we always have
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa).$
@@
Recall that $f,g$ are both weakly $\kappa$-bi-Lipschitz. Applying Lemma 2.15 and Lemma 2.18, one can verify that
@@
$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa)$
$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa).$ (29)
@@
Note that $D_{n}(Z)$, $D_{n}(Y^{\prime})$ are conditionally i.i.d. given $D_{n}(X,X^{\prime})$. Also, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime\prime})$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$. By Lemmas 2.8 and 2.19,
@@
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq 2H_{n}(f(X)+g(Y))+2H_{n}(X)+2H_{n}(Y)-2H_{n}(X,Y)+O(\log\kappa).$
@@
$\qquad\qquad+2\big{(}H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big{)}+H_{n}(X)+H_{n}(Y)+O(\log\kappa).$
$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa).$
We are ready to present a proof of Theorem 3.1
@@
$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa)$
@@
$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+\frac{H_{n}(X)+H_{n}(Y)}{2}+\log K_{2}$
$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa).$ (38)
@@
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa).$
@@
4.1 Diagonal forms with independent random variables
In this section, we consider the diagonal real quadratic form $\phi(x,y)=\alpha x^{2}+\beta y^{2}$. When we write $O_{\phi}(\cdot)$, it means that the implied constant may depend on the coefficients $\alpha$ and $\beta$ that occurred in $\phi$.

### v4

@@ 2.4 Frostman measures
(iii) We say a random variable $X$ is conditionally $(s,C)$-Frostman given $Y$, if
@@
for any ball $B(x,r)$ and Borel set $E$. Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
The following properties follow directly from the definitions, whose proofs are left as an exercise.
@@
(iv) Suppose that $(X,Y)$ is a jointly $(s_{1},s_{2},C)$-Frostman random variable whose range is contained in $B(0,R)$. Then $(X,Y)$ is marginally $(s_{1},CR^{s_{2}};s_{2},CR^{s_{1}})$-Frostman.
@@
Assume that $b\neq 0$. If $q_{i}>0$, then $Y$ is located in an interval $J_{i}$ of length at most $\frac{(|a|+2)r}{|b|}$. Hence,
@@
First, consider the case $ad\neq 0$. When $r_{1}\leq r_{2}$, let $\{J_{i}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r_{1}$. We estimate $p$ by
@@
One can check that, if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{(|b|+2)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
@@
We apply similar arguments as above. For non-zero terms, the value of $Y$ locates in an interval $J_{i}^{\prime}$ of length at most $\frac{(|c|+2)r_{2}}{|d|}$. Therefore,
@@
(i) The conclusion comes from [12, Lemma 4.8].
@@
Let $m\geq 1$. Suppose that $Z$ is a random variable with bounded range in ${\mathbb{R}}^{m}$. Assume that there are numbers $s,C>0$ such that $P(|Z|<r)\leq Cr^{s}$ for any $r>0$. Let $0<\eta\leq\min\{2^{-1}(2C)^{-1/s},\,1/2\}$, and denote by $E$ the event that $|Z|\geq\eta$. Assume that $P(E)\gg 1$. Then the random variable $Z_{E}$ satisfies
$H_{n}(Z_{E})=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
@@
Recall that $Z_{E}$ is the random variable such that $Z_{E}\equiv(Z|E)$, and $P_{E}$ is the conditional probability measure given $E$ (see Lemma A.2).
Assume that $k_{0}$ is the natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$. Combining the conditions in this lemma, one deduces that
$H_{n}(Z_{E})$ $=-\sum\limits_{k\in\mathbb{Z}^{m}\atop|k|\geq k_{0}+m}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})$
$\qquad+O\left(-\sum\limits_{k_{0}-m\leq|k|<k_{0}+m}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})\right)$
$=-\sum\limits_{|k|\geq k_{0}+m}\frac{P(D_{n}(Z)=k2^{-n})}{P(E)}\big(\log P(D_{n}(Z)=k2^{-n})-\log P(E)\big)+O_{m}(1)$
$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+m}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})+O_{m}(1).$
First, we analyze the case that $\eta>m2^{-n+2}$, which gives $(k_{0}+2m)2^{-n}\leq\eta+(2m+1)2^{-n}<2\eta$. Applying Lemma 2.9 with the fact
$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(Z)=k2^{-n})\leq P(|Z|<(k_{0}+2m)2^{-n})\leq P(|Z|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
@@
$H(Z_{E})$ $=\frac{1}{P(E)}\left(H_{n}(Z)+\sum\limits_{|k|<k_{0}+m}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})\right)+O_{m}(1)$
$=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
@@ 2.5 Entropy of the distance function
Let $\mu$ be a probability measure with $\operatorname{spt}(\mu)\subseteq B(0,R)\subseteq\mathbb{R}^{2}$, for some $R>0$. Then, for any $0<t\leq 1$ and for real numbers $0<\eta<r<2R$, we have $\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)$.
@@
Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subseteq B(0,R)\subseteq\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$ and $C,R>0$. Then for any $r,\eta,t$ with $0<\eta<r<2R$, $0<t\leq 1$ and $t<2s-\frac{1}{2}$, it satisfies that
@@
$\qquad=O_{t}\left(\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}\right)+O(\tilde{R}^{-t-1/2}):=M.$
@@
$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O(\eta^{t}).$
@@
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subseteq\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and s<t<2s1/2s<t<2s-1/2, we have
@@
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subseteq\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
@@
Let $\eta$ be a parameter with $0<\eta<\min\{2^{-1}(2C)^{-1/2s},\,1/2\}$ to be determined later. Denote by $E$ the event that $|Z|\geq\eta$. By Lemma 2.34, one has $P(|Z|<\eta)\leq C\eta^{2s}$, which gives $P(E)\geq 1/2$. By Lemma 2.33, we deduce that
@@
$H_{n}(Z^{2})$ $\geq P(E)\,H_{n}(Z_{E}^{2})\geq P(E)\left(H_{n}(Z_{E})+O\left(\log\frac{1}{\eta}\right)\right)$
@@
Verification shows that the requirements involving $t$ and $\eta$ are satisfied, and $2s-t-1/2\gg_{s}1$. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
@@
Let $X$, $X^{\prime}$, $Y$, $Y^{\prime}$ be four independent real random variables. Suppose that $X$ and $X^{\prime}$ are both $(s,C)$-Frostman with ranges contained in [c,c][-c,c], where $0<s\leq 1$ and $C,c>0$. Then, for any $n\in\mathbb{N}$,
@@
By Lemmas 2.30 and 2.31, the random variable $Z:=X-X^{\prime}$ is $(s,O(C))$-Frostman. By Lemma 2.32, one has $H_{n}(Z)\geq ns-\log C+O(1)$. Also, $range(Z)\subseteq[-2c,2c]$. Let $E$ be the event that $|Z|\geq\eta$ as in Lemma 2.33. Follow the same arguments as in the proof of Theorem 2.41, we obtain that
@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
In this section, we follow the method introduced by Tao in [26, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem. We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section.
@@
Let $(X,Y)$ be a random variable, and $n$ be a natural number. Suppose that
@@
In addition, for any real valued functions $f$ and $g$ that are weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(X)$ and $range_{+}(Y)$, respectively, with
@@
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$ (24)
Here $\log K_{1},\,\log K_{2},\,\log\kappa\,\gg 1$ and $d\in\mathbb{N}$.
@@
Let $(X,Y)$ be a random variable. Then for any $n\in\mathbb{N}$, there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that (i)–(iv) in Theorem 3.1 hold, with $X^{\prime\prime}$ and $Y^{\prime\prime}$ being conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$.
@@
Let $(X,Y)$ be a random variable, and $n$ be a natural number. Let $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ be the random variable shown in Lemma 3.3. For any function $h$ that is $\kappa$-Lipschitz on $range_{+}((X,Y))$, the following properties hold.
@@
Noting that $(Y^{\prime},D_{n}(X^{\prime}))\equiv(Y^{\prime\prime},D_{n}(X^{\prime}))$, so
@@
Combining these with (i), one gets (ii). Claim (iii) follows by a similar argument. ∎
@@
Under the notations and conditions stated in Theorem 3.1, we have
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d).$
@@
Recall that $f,g$ are both weakly $(\kappa,d)$-bi-Lipschitz. Applying Lemma 2.16 and Lemma 2.18, one can verify that
@@
$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa)+O(\log d)$
$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa)+O(\log d).$ (25)
@@
Note that $D_{n}(Z)$, $D_{n}(Y^{\prime})$ are conditionally i.i.d. given $D_{n}(X,X^{\prime})$. Also, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime\prime})$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$. By Lemma 2.8 and Corollary 3.4,
@@
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})$
$\qquad\leq 2H_{n}(f(X)+g(Y))+2H_{n}(X)+2H_{n}(Y)-2H_{n}(X,Y)+O(\log\kappa)+O(\log d).$
@@
$\qquad\qquad+2\big(H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big)+H_{n}(X)+H_{n}(Y)+O(\log\kappa)+O(\log d).$
$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d).$
We now present the proof of Theorem 3.1.
@@
$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa)+O(\log d)$
@@
$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d)$ (34)
$\qquad\qquad\qquad+\frac{H_{n}(X)+H_{n}(Y)}{2}+\log K_{2}$
$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$ (35)
@@
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$
@@
4.1 Diagonal quadratics with independent random variables
In this subsection, we first consider the diagonal real quadratic form $\phi(x,y)=\alpha x^{2}+\beta y^{2}$.